### v2

# zkCraft: Prompt-Guided LLM as a Zero-Shot Mutation Pattern Oracle for TCCT-Powered ZK Fuzzing

Abstract
Zero-knowledge circuits enable privacy-preserving and scalable systems but are difficult to implement correctly due to the tight coupling between witness computation and circuit constraints. We present zkCraft, a practical framework that combines deterministic, R1CS-aware localization with proof-bearing search to detect semantic inconsistencies. zkCraft encodes candidate constraint edits into a single Row-Vortex polynomial and replaces repeated solver queries with a Violation IOP that certifies the existence of edits together with a succinct proof. Deterministic LLM-driven mutation templates bias exploration toward edge cases while preserving auditable algebraic verification. Evaluation on real Circom code shows that proof-bearing localization detects diverse under- and over-constrained faults with low false positives and reduces costly solver interaction. Our approach bridges formal verification and automated debugging, offering a scalable path for robust ZK circuit development.
Keywords zero-knowledge proofs, fuzzing, Circom, TCCT, Row-Vortex commitment, Violation IOP, large language models, mutation testing
1 Introduction
Zero-knowledge circuits provide succinct, verifiable proofs while exposing minimal information about private inputs, and they form the foundation of many privacy and scalability systems. Writing correct circuits is difficult because the witness generator that computes private values must align precisely with the constraint system used for proof verification. When this alignment fails, two broad classes of faults appear. Under-constrained faults arise when the constraint system admits execution traces that no valid witness can produce. Over-constrained faults arise when a valid witness-producing execution is rejected by the constraints. Both fault classes have important security and operational consequences in deployed systems and have driven significant recent research efforts  [41, 55, 14, 56, 44, 34].
Prior work addresses these problems from multiple perspectives. Static and pattern-based detectors reveal common syntactic mistakes but often miss semantic discrepancies that manifest only during witness generation or under specific inputs  [13, 53, 49]. Formal verification and SMT-based analyses can provide strong guarantees but typically do not scale to large circuits, and they struggle with the combined search space of program edits and input assignments  [54, 47, 24]. Dynamic approaches based on fuzzing and metamorphic testing improve practical bug-finding coverage, yet they frequently depend on many expensive solver calls or unguided mutations that produce numerous low-value candidates  [41, 55, 56]. Recent work on verifier-side encodings, interactive oracle proofs, and efficient polynomial commitments offers promising primitives to compress or certify search outcomes  [44, 64, 65]. In parallel, advances in model-guided mutation and LLM-driven fuzzing have shown that language models can generate semantically rich test patterns and seed mutations for security testing  [37, 62, 33, 63, 4, 29].
Despite these developments, important gaps remain. Most algebraic and IOP-style techniques have not been integrated with practical program-level mutation strategies, so solver costs remain a bottleneck for end-to-end bug discovery. Conversely, LLM-guided mutation methods improve suggestion quality but they are typically detached from proof-bearing search and therefore can produce proposals that are expensive to validate or that lead to many false positives  [50, 29, 62]. Compilers and zkVM toolchains introduce additional complexity related to optimization and backend behavior, which further complicates end-to-end detection and increases the cost of exhaustive solver-based validation  [18, 59, 66, 57].
Motivated by these observations, we design zkCraft, a ZK-native pipeline that tightly couples algebraic localization, succinct violation proofs, and deterministic, prompt-guided mutation templates. zkCraft reduces the combinatorial task of finding small, vulnerability-inducing edits to a single algebraic existence statement. It constructs compact per-row fingerprints to prioritize a limited candidate pool, encodes row selection and substitution parameters as a Row-Vortex polynomial, and asks a prover to produce a Violation interactive oracle proof that certifies the existence of a witness satisfying the edited constraints while yielding a different public output. The proof serves as an auditable counterexample from which concrete substitutions can be recovered. To focus search on high-yield edits, zkCraft employs deterministic LLM-driven mutation templates as a zero-shot pattern oracle for right-hand-side edits and for input-sampling heuristics. These templates propose algebraic edits but do not replace the proof-producing components.
Our contributions are as follows. We present a ZK-native search framework that maps the search for small, vulnerability-causing edits to a single algebraic existence statement encoded as a Row-Vortex polynomial and certified with a Violation IOP. We design deterministic LLM interfaces that act as zero-shot mutation-template oracles to propose right-hand-side edits and input-sampling biases while remaining decoupled from proof generation. We describe practical implementation techniques for compact per-row fingerprinting, a scoring heuristic to select a small candidate pool, and a deterministic finite-field synthesis that recovers concrete substitutions from proof-embedded encodings. We evaluate the approach on representative Circom circuits and zkVM workloads and show that the ZK-native workflow discovers diverse under- and over-constrained faults while substantially reducing solver calls and producing auditable counterexamples suitable for developer inspection.
2 Related Work
2.1 Correctness and verification of ZK pipelines and compilers
The correctness of zero-knowledge compilation and circuit-processing pipelines has attracted growing attention because compiler or pipeline bugs can yield accepted yet invalid proofs. Static and dynamic analyses specifically targeting ZK circuit representations and their compilers have been proposed to reveal semantic and logic flaws early in the toolchain. Prior work presents static detectors over circuit dependence graphs and evaluates them on large Circom codebases, demonstrating that semantic checks can recover subtle vulnerabilities in real projects  [55]. Metamorphic and mutation-driven approaches have been adapted to the ZK compiler setting to detect incorrect compilation behaviors by generating semantically equivalent variants and checking for inconsistent outputs  [56, 28]. Systematic fuzzing of end-to-end ZK pipelines has also been shown to surface logic bugs that elude conventional testing, illustrating the importance of domain-aware test oracles for pipelines that span DSL, compiler, and proof-generation layers  [22, 10].
2.2 Fuzzing, metamorphic testing, and oracles for ZK systems
Fuzzing and metamorphic testing have been successfully adapted to find pipeline-level and finalization bugs in ZK deployments, including rollup finalization failures and pipeline logic flaws; these techniques emphasize careful oracle design and input-generation strategies to trigger diverse behaviors  [28, 43]. Tools that combine metamorphic relations with fault-injection or targeted mutation expose both soundness and completeness weaknesses in zkVMs and related infrastructures  [23, 28]. Work that specifically targets the oracle problem in ZK fuzzing highlights the need for proofs or semantic checks that distinguish genuine counterexamples from benign variations  [10, 22]. Complementary studies show that bounded, structured mutation templates and algebraic checks outperform unguided random mutation when the objective is to find under-constrained or mis-specified constraints  [15, 25].
2.3 LLMs and automated test/mutation generation
Large language models have been explored as code-generation and test-generation oracles across multiple domains. In compiler and test-generation settings, LLMs can be guided to produce high-quality test programs or mutation patterns when fed targeted prompts and feedback, enabling discovery of deep optimization and correctness bugs  [61, 35, 12]. Recent work combines LLMs with mutation testing and iterative refinement to generate tests that kill mutants and to produce sampling strategies tailored to observed failures  [21, 20, 48]. Studies that evaluate LLMs specifically for ZK program generation report that off-the-shelf models often produce syntactically plausible but semantically incorrect ZK code, motivating augmentation strategies such as constraint sketching and guided retrieval  [60, 27]. These findings motivate zkCraft’s design choice to treat LLMs as deterministic template oracles rather than as formal solvers  [60, 27].
2.4 Symbolic execution, SMT guidance, and hybrid strategies
Symbolic execution enhanced with solver-time prediction and synthesized solver strategies has been shown to improve scalability for complex programs, and these ideas inform hybrid pipelines that combine symbolic techniques with search or sampling  [35, 12, 52]. For ZK circuit mutation and verification, solver-guided fallback channels that ground small template families and use finite-field SMT/SAT queries remain practical when pure ZK-native IOPs are infeasible; this hybrid approach balances completeness and engineering constraints  [5, 2]. Machine-checked proofs and formally verified protocol implementations provide strong guarantees and serve as an orthogonal means to reduce the trusted code base for ZK constructions  [2, 5].
2.5 ZK-native proof techniques, commitments, and succinct verification
Recent advances in polynomial commitments, IOPs and succinct proof systems enable encoding higher-level search oracles as compact ZK statements. Systems that adopt ZK-native encodings show how commitment-and-IOP primitives can be repurposed to certify existential properties about edits and witnesses, thereby converting existential search into a single verifiable artifact  [38, 6, 3]. GPU-accelerated proof systems and hardware-algorithm co-design work illustrate routes to practical prover performance that scale to larger circuits  [36, 9, 46]. Benchmarks and tooling that provide comparative performance baselines for SNARK stacks are useful for choosing commit/IOP backends when building ZK-native workflows  [15, 32].
2.6 Circuit languages, compilers and verification frameworks
Domain-specific languages for ZK, their compiler frontends, and associated verification toolchains form a rich ecosystem. Efforts to improve circuit description languages and make them amenable to static verification encourage safer ZK application development  [6, 40]. Compiler-focused fuzzing and white-box analysis have uncovered deep logic bugs in real-world compilers, motivating tighter integration of source-level diagnostics with pipeline-level checks  [61, 56]. Tools that certify circuits with refinement types or other higher-level specifications reduce a class of errors by construction and complement testing-based approaches  [31, 25].
2.7 Applications, surveys, and domain-specific ZK use cases
Zero-knowledge proofs are being applied widely, from verifiable decentralized machine learning to privacy-preserving credentialing and domain-specific verification tasks  [58, 7, 17]. Surveys and systematizations contextualize verifiability challenges and catalog practical design patterns and open problems for ZKP-based applications  [26, 30]. Domain work that integrates anomaly detection, sensing, or constrained hardware contexts highlights practical verification needs in applied settings  [45, 42, 1].
2.8 Mutation tooling, scientific debugging and cross-domain testing
Mutation testing and scientific debugging have been adapted to multiple domains including quantum circuits and hardware verification; these adaptations emphasize careful mutant design and oracle selection to obtain meaningful faults  [19, 39]. Iterative, model-guided debugging pipelines that combine LLMs with mutation testing demonstrate that principled human-in-the-loop strategies and automated refinement can improve fault detection while keeping developer review effort manageable  [21, 20, 48]. These cross-domain lessons guided the design of zkCraft’s mutation templates and deterministic post-processing to keep generated edits syntactically valid and semantically focused  [60, 21].
2.9 Summary of positioning
In summary, prior work supplies three families of building blocks for zkCraft. The first family covers ZK pipeline correctness, compilers, and static/dynamic detectors that expose semantic vulnerabilities  [55, 56, 22, 28]. The second family comprises mutation, fuzzing, and LLM-guided test generation techniques that provide high-yield input and edit proposals  [61, 21, 20, 48]. The third family contains ZK-native proof primitives, commitment/IOP engineering, and performance tooling that enable replacing many solver calls with compact cryptographic artifacts  [38, 3, 36, 32]. zkCraft integrates elements from all three families by combining algebraic localization, a Row-Vortex commitment with a Violation IOP, and deterministic LLM-driven mutation templates to produce proof-bearing counterexamples that are both practical and auditable  [51, 25, 27]. guarantees, advancing the state-of-the-art in ZKP security.
Figure 1: Overview of the zkCraft framework for ZK-native fuzzing and mutation. The pipeline begins at the Input Layer, where a Circom program is decomposed into R1CS matrices. In Stage 1 (The Slicer), a Sparse Fingerprint Scanner computes diagnostic scores to prune the constraint space into a manageable Candidate Pool $\mathcal{R}_{\mathrm{cand}}$. This process is accelerated by the LLM Oracles, which provide zero-shot Mutation Templates and learn vulnerability patterns to bias the search. In Stage 2 (ZK-Native Engine), candidate edits are bundled into the Row-Vortex Commitment $R(X,Y)$. A Violation IOP (utilizing algebraic Sum-Check identities) then searches for a witness that satisfies the edited constraints while diverging from the original output. The resulting Succinct Proof $\pi$ serves as a Proof-as-Counterexample, from which the Witness Extraction machinery reconstructs the concrete trace to confirm the vulnerability via TCCT.
Figure 2: zkCraft versus baseline tools. The horizontal axis uses a logarithmic scale from one second to two hours. Curves show cumulative unique true positives averaged over five seeds and shaded bands indicate one standard deviation. A gold marker at 100 seconds highlights early convergence.
Figure 3: Ablation results. Bars represent the log10 of the slowdown factor relative to the default configuration. The red line reports the number of bugs not discovered within 24 hours when a single module is removed. A horizontal threshold marks a 100 times slowdown.
Figure 4: zkCraft convergence under four mutation/population settings. Time (log scale) on the horizontal axis; cumulative true positives on the vertical. Gold dashed line marks 100-second convergence; shaded areas show one standard deviation.
Figure 5: Row-Vortex commitment schematic. The three panels illustrate candidate row selection, construction of the bivariate polynomial, and the Violation IOP that yields a succinct proof. The top annotation summarizes the replacement of many SMT calls by a single commitment.
Figure 6: Proof size vs. constraint count (top) and prover time (bottom), both on log-scaled axes. Dashed line shows proportional growth; horizontal marker highlights small proof size in ZK-native mode.
Figure 7: Representative confirmed vulnerabilities and corresponding counterexamples. Left panels show compact code excerpts with highlighted weak assignments. Right panels list concrete input values and the observed outputs that demonstrate the fault.
Figure 8: Strategy comparison as a radar plot. Axes report early discovery at two short checkpoints, final discovery at two hours, false positive tendency after inversion, and normalized median runtime after inversion. Polygons show the mean over five seeds.
Figure 9: Hyperparameter sensitivity heatmap. The horizontal axis is mutation strength and the vertical axis is population size. Colors encode the number of true positives found within the two-hour budget. The default configuration is framed and annotated.
Figure 10: Sensitivity of zkCraft to key hyperparameters across four settings. Each panel varies one parameter with others fixed; curves show cumulative true positives over time (log scale), averaged across five seeds. A gold dashed line at 100 seconds marks the early-convergence regime.
3 Methodology
This section formalizes the zkCraft methodology and its ZK-native extension. We present a deterministic, R1CS-aware pipeline that replaces randomized program mutation with a small-scope, algebraic slicing procedure. The novel extension replaces the external SMT/SAT oracle with a succinct zero-knowledge protocol. Under this design the existence of a small edit set and the concrete edits themselves are certified by a single proof $\pi$; proof verification both confirms the vulnerability and yields the concrete counterexample trace without a separate TCCT [51] invocation.
3.1 Notation and preliminaries
Let $\mathbb{F}_{q}$ denote a prime finite field of order $q$. An R1CS instance with $m$ constraints and $n$ variables is represented by matrices $A,B,C\in\mathbb{F}_{q}^{m\times n}$. For each row index $i\in\{1,\dots,m\}$, let $a_{i},b_{i},c_{i}$ denote the $i$-th rows of $A,B,C$, respectively. A witness vector is $w\in\mathbb{F}_{q}^{n}$.
$\langle a_{i},w\rangle\cdot\langle b_{i},w\rangle\;=\;\langle c_{i},w\rangle,$ (1)
where $\langle\cdot,\cdot\rangle$ denotes the inner product over $\mathbb{F}_{q}$ and Eq. (1) expresses the $i$-th Rank-1 constraint.
$F(w)\;=\;\big(\langle a_{i},w\rangle\cdot\langle b_{i},w\rangle-\langle c_{i},w\rangle\big)_{i=1}^{m}.$ (2)
where $F(w)$ is the vector of per-row residuals and a valid witness satisfies $F(w)=0$.
We use the Trace-Constraint Consistency Test (TCCT) as semantic acceptance: a mutated program $P^{\prime}$ and input $x^{\prime}$ produce a TCCT counterexample if the execution trace $(z^{\prime},y^{\prime})$ satisfies the original constraints $C$ while yielding public output $y^{\prime}\neq y$, where $y$ denotes the original public output for the relevant input.
3.2 Formal problem statement
Given an R1CS $\mathcal{R}=(A,B,C)$ over $\mathbb{F}_{q}$, a program $P$ that produces a witness $w$ on input $x$, a mapping from source-level weak-assignment sites to R1CS rows, and an edit budget $t_{\max}\in\mathbb{N}$, we seek a small collection of right-hand-side edits (restricted to weak assignments) producing $P^{\prime}$ and a witness $w^{\prime}$ for some input $x^{\prime}$ such that
$P^{\prime}(x^{\prime})=(z^{\prime},y^{\prime}),\qquad C(x^{\prime},z^{\prime},y^{\prime})=\textsf{true},\qquad y^{\prime}\neq y.$ (3)
where $y$ denotes the public output of the original program on the considered input. The decision variant asks whether such a set of edits of size at most $t_{\max}$ exists. Because TCCT decision problems are hard in the worst case, we optimize for small edit cardinalities via a focused search.
3.3 Stage 1: Sparse signature extraction
For every constraint row $i$ compute the nonzero index supports
$\mathrm{supp}(a_{i}),\ \mathrm{supp}(b_{i}),\ \mathrm{supp}(c_{i}),$ (4)
where $\mathrm{supp}(\cdot)$ returns the set of indices with nonzero coefficients.
Define two integer diagnostics per row:
$\kappa^{\mathrm{w}}_{i}$ $:=\big|\big(\mathrm{supp}(a_{i})\cup\mathrm{supp}(b_{i})\cup\mathrm{supp}(c_{i})\big)\cap\mathcal{W}\big|,$ (5)
$\kappa^{\mathrm{c}}_{i}$ $:=\big|\big(\mathrm{supp}(a_{i})\cup\mathrm{supp}(b_{i})\cup\mathrm{supp}(c_{i})\big)\cap\mathcal{K}\big|,$ (6)
where $\mathcal{W}$ indexes witness/intermediate variables and $\mathcal{K}$ indexes constant/public/simple-expression terms.
Form a compact per-row fingerprint and a scalar score:
$\mathrm{fp}_{i}\;=\;\mathrm{Trunc64}\big(\mathrm{Hash}(\mathrm{coeffs}(a_{i},b_{i},c_{i})\,\|\,\kappa^{\mathrm{w}}_{i}\,\|\,\kappa^{\mathrm{c}}_{i})\big),$ (7)
where $\mathrm{Hash}$ is a collision-resistant hash and $\mathrm{Trunc64}$ truncates to 64 bits; and
$s_{i}\;=\;\lambda\cdot\frac{\kappa^{\mathrm{c}}_{i}}{\kappa^{\mathrm{w}}_{i}+1}\;-\;\mu\cdot\kappa^{\mathrm{w}}_{i},$ (8)
where $\lambda,\mu>0$ are tunable scalars and the fractional term stabilizes ratio sensitivity. Select the top-$k$ rows with smallest $s_{i}$ to form the candidate pool $\mathcal{R}_{\mathrm{cand}}$ (typical $k\leq 32$). This stage runs in time linear in the number of nonzeros.
3.4 Stage 2: Row-Vortex commitment and Violation IOP (ZK-native)
Row-Vortex is a bivariate polynomial encoding that combines candidate row selections and substitution constants for R1CS constraints, providing a compact representation of mutation choices within the zkCraft framework. We replace the loop of SMT/SAT queries by a single zero-knowledge interactive proof that simultaneously attests to the existence of a small set of edits and carries a compact counterexample when one exists. The protocol works on a preselected candidate set $\mathcal{R}$ and encodes both binary row selection and per-site substitutions into a two-variable polynomial, which we call the Row-Vortex polynomial:
$R(X,Y)$ $=\sum_{i\in\mathcal{R}}\delta_{i}\,\mathrm{row}_{i}(X)\;+\;c_{i}\,\mathrm{sel}_{i}(Y).$ (9)
where $\delta_{i}\in\{0,1\}$ denotes whether row $i$ is chosen, $c_{i}\in\mathbb{F}_{q}$ is the substitution constant associated with row $i$, $\mathrm{row}_{i}(X)$ is a low-degree univariate encoding of the $i$-th R1CS row in variable $X$, $\mathrm{sel}_{i}(Y)$ isolates the $i$-th slot via $Y$, and $\mathcal{R}$ is the candidate index set.
The prover commits to $R(X,Y)$ under an efficient polynomial commitment scheme (for example, contemporary HyperPlonk+/BaseFold-style constructions). The commitment $\mathsf{Comm}(R)$ binds the prover to a unique pair $(\delta,c)$ while remaining succinct. The prover additionally holds a candidate witness $w^{\prime}$ that, together with $(\delta,c)$, is intended to satisfy the edited constraint system.
To express the public-output divergence in algebraic form, fix distinct evaluation points $\{u_{j}\}_{j=0}^{m_{\mathrm{out}}-1}$ for the public-output coordinates and let $L_{j}(X)$ be the corresponding Lagrange basis polynomials. Define the multilinear output-difference polynomial
$\Delta_{\mathrm{out}}(X)\;=\;\sum_{j=0}^{m_{\mathrm{out}}-1}\big(y^{\prime}_{j}-y_{j}\big)\,L_{j}(X).$ (10)
where $y^{\prime}_{j}$ denotes the $j$-th coordinate of the public output produced by $w^{\prime}$, $y_{j}$ denotes the original public output coordinate, and $L_{j}(X)$ is the degree $\leq m_{\mathrm{out}}-1$ Lagrange basis polynomial for point $u_{j}$. The predicate $y^{\prime}\neq y$ is equivalent to the polynomial inequality $\Delta_{\mathrm{out}}(X)\not\equiv 0$.
Aggregate the edited constraint residuals into a single per-evaluation polynomial by evaluating the Row-Vortex encoding on an evaluation domain $\mathcal{D}$ (for example a subset of $\{0,1\}^{\ell}$) and summing the pointwise residual encodings:
$\Phi_{R}(U)\;=\;\sum_{i\in\mathcal{D}}\phi_{i}\big(R(U,\cdot),\,w^{\prime}\big).$ (11)
where $\mathcal{D}$ is the chosen evaluation domain, $\phi_{i}(\cdot)$ denotes the algebraic form of the $i$-th constraint residual when evaluated at domain point $U$, and $R(U,\cdot)$ is the polynomial in the remaining variable obtained by fixing the first argument $X=U$.
We capture the localization predicate, namely the existence of edits that zero the edited residuals while changing the public output, by the following existential algebraic statement:
$\exists\;($ $\delta\in\{0,1\}^{|\mathcal{R}|},\;c\in\mathbb{F}_{q}^{|\mathcal{R}|},\;w^{\prime}\in\mathbb{F}_{q}^{n})\quad\text{s.t.}$ (12)
$F_{\delta,c}(w^{\prime})=0\quad\text{and}\quad\Delta_{\mathrm{out}}(X)\not\equiv 0.$ (13)
where $F_{\delta,c}:\mathbb{F}_{q}^{n}\to\mathbb{F}_{q}^{m}$ denotes the vector of edited per-row residuals obtained after applying the selection $\delta$ and substitutions $c$; $F_{\delta,c}(w^{\prime})=0$ means the edited R1CS is satisfied by $w^{\prime}$, and $\Delta_{\mathrm{out}}(X)\not\equiv 0$ arithmetizes the requirement that the public outputs differ.
The prover reduces the existential statement (13) to a Sum-Check style identity over $\mathcal{D}$. Concretely, form the global algebraic sum
$\sum_{U\in\mathcal{D}}\Phi_{R}(U)\,\Delta_{\mathrm{out}}(U)\;=\;0.$ (14)
where the integrand pairs the edited residual aggregation $\Phi_{R}(U)$ with the arithmeticized divergence $\Delta_{\mathrm{out}}(U)$, so that a valid triple $(\delta,c,w^{\prime})$ makes the summand vanish pointwise and the global sum evaluate to zero.
The prover runs a standard public-coin Sum-Check/IOP over the committed polynomial. Let $t=\lceil\log_{2}|\mathcal{D}|\rceil$ and decompose $\mathcal{D}=\mathcal{D}_{1}\times\cdots\times\mathcal{D}_{t}$. The Sum-Check transcript is produced as a sequence of univariate polynomials $g_{1},\dots,g_{t}$ satisfying the iterated-sum relations:
$\sum_{U\in\mathcal{D}}\Phi_{R}(U)\,\Delta_{\mathrm{out}}(U)$ $=\sum_{u_{1}\in\mathcal{D}_{1}}g_{1}(u_{1}),$ (15)
$g_{k}(u_{1},\dots,u_{k})$ $=\sum_{u_{k+1}\in\mathcal{D}_{k+1}}g_{k+1}(u_{1},\dots,u_{k+1})$ (16)
$\quad\text{for }k=1,\dots,t-1.$ (17)
where each $g_{k}$ is univariate in its last argument (degree bounded by a small constant determined by the encoding), and the equalities enforce consistency between the iterated sums and the claimed global value.
During the interactive protocol the verifier samples independent random challenges $\zeta_{1},\dots,\zeta_{t}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{F}_{q}$ and asks the prover to open committed objects at these points. Concretely the verifier requests evaluations of
$R(\zeta_{k},\,Y),\qquad\Phi_{R}(\zeta_{k}),\qquad\Delta_{\mathrm{out}}(\zeta_{k})$ (18)
for the appropriate indices $k$. Here $R(\zeta_{k},Y)$ denotes the univariate polynomial in $Y$ obtained by substituting $X=\zeta_{k}$, $\Phi_{R}(\zeta_{k})$ is the aggregated residual value at $U=\zeta_{k}$, and $\Delta_{\mathrm{out}}(\zeta_{k})$ is the divergence test evaluation at the same point.
In the final step, the prover opens the commitment $\mathsf{Comm}(R)$ at the sampled $X$-coordinates and supplies evaluation proofs. The verifier checks that the openings are consistent with $\mathsf{Comm}(R)$ under the chosen commitment scheme, that the Sum-Check consistency relations hold at the sampled challenges, and that the divergence indicator is nonzero for at least one sampled challenge, namely $\Delta_{\mathrm{out}}(\zeta_{k})\neq 0$ for some $k$. The verifier accepts exactly when all checks succeed.
Soundness follows from two mechanisms. The Sum-Check subprotocol forces the prover to respect the global algebraic identity except with negligible probability, and randomized evaluation of $\Delta_{\mathrm{out}}$ yields a Schwartz–Zippel-style guarantee that a nonzero polynomial will be detected with high probability when the field size $q$ is large relative to $m_{\mathrm{out}}$. Upon acceptance, an extractor operating under standard commitment/IOP assumptions can recover the committed low-degree encodings and reconstruct the finite tuple $(\delta,c,w^{\prime})$ or directly obtain the concrete counterexample $(x^{\prime},z^{\prime},y^{\prime})$; thus the accepted transcript functions both as a succinct certificate of existence and as a carrier of the execution-level counterexample for developer triage. As detailed in Subsection 6.7, the degree bounds and node selection ensure injective encoding and stable extraction.
3.5 Proof-as-counterexample (verification and extraction)
A verifier’s successful validation of the succinct proof $\pi$ certifies the existence of a witness satisfying the target algebraic statement in Eq. (13). Our practical extraction policy is two-fold: either the verifier reconstructs the finite tuple $(\delta,c,w^{\prime})$ from openings produced against the polynomial commitment $\mathsf{Comm}(R)$, or the proof itself contains an auditable encoding of the concrete counterexample $(x^{\prime},z^{\prime},y^{\prime})$. In both cases, verification not only attests to the bug but also yields the concrete execution trace without necessitating an independent TCCT replay; in effect, the proof functions as the counterexample certificate. Where necessary for developer inspection, the verifier may still optionally re-run the mutated program for richer runtime traces.
From commitment opening to counterexample recovery
We now present a deterministic procedure the verifier (or an extractor acting on behalf of the verifier) executes after accepting an IOP transcript. The procedure consumes the opened evaluations and their evaluation proofs and reconstructs the compact data $(\delta,c,w^{\prime})$; when the IOP transcript already carries an encoding of the witness polynomial $W^{\prime}(X)$ the procedure is simpler because interpolation yields $w^{\prime}$ directly.
1
Input : Accepted transcript containing commitment openings and evaluations, and evaluation proofs
Output : Recovered tuple $(\delta,c,w^{\prime})$ and the synthesized program $P^{\prime}$ (or fail)
2 Collect opened values and their proofs: $\{R(\zeta_{k},Y),\Phi_{R}(\zeta_{k}),\Delta_{\mathrm{out}}(\zeta_{k}),\pi_{\mathrm{eval}}\}_{k}$
3 Compute $\rho$, the compact vector derived from the opened encodings (aggregation step)
4 Compute $(\delta,c)\leftarrow M^{-1}\cdot\rho$ using the inverse affine map associated with the encoding
5 if witness polynomial $W^{\prime}(X)$ is present in the transcript then
6   Interpolate $W^{\prime}(X)$ to obtain the vector $w^{\prime}$;
7 else
8   Form the algebraic reconstruction system $F_{\delta,c}(w^{\prime})=0$ derived from the substituted constraints
9   Solve for $w^{\prime}$ by Gaussian elimination or an $n\times n$ linear solve over $\mathbb{F}_{q}$;
10 Synthesize mutated source $P^{\prime}$ by substituting the recovered constants $c$ into the identified weak-assignment sites
11 Derive public inputs $x^{\prime}$ from the public entries of $w^{\prime}$ and output the concrete counterexample $(x^{\prime},z^{\prime},y^{\prime})$
12 return $(\delta,c,w^{\prime})$
Algorithm 1 Extractor: recover concrete counterexample from an accepted IOP transcript
We next state cost estimates for this deterministic extraction. Let $k=|\mathcal{R}|$ denote the number of candidate rows encoded in the Row-Vortex polynomial and let $n$ denote the witness dimension. Recovering the compact selection vector $(\delta,c)$ from the opened encodings requires applying the inverse affine map $M^{-1}$. When implemented with a sparse reconstruction routine this step runs in time
$T_{\mathrm{select}}\;=\;O(k\log k),$ (19)
where $k$ is the size of the candidate set and the logarithmic factor accounts for sorting/hashed-sparse recovery primitives.
If the transcript contains an encoded witness polynomial $W^{\prime}(X)$, interpolation recovers $w^{\prime}$ in time
$T_{\mathrm{interp}}\;=\;O(n\log n),$ (20)
where $n$ is the number of witness coordinates and the complexity assumes FFT-style interpolation over a suitable evaluation domain.
If $W^{\prime}(X)$ is not provided, the extractor assembles the linear/algebraic reconstruction system
$F_{\delta,c}(w^{\prime})\;=\;0,$ (21)
where $F_{\delta,c}:\mathbb{F}_{q}^{n}\to\mathbb{F}_{q}^{m}$ denotes the vector of edited residuals after applying $(\delta,c)$. Solving Eq. (21) by standard Gaussian elimination costs
$T_{\mathrm{solve}}\;=\;O(n^{\omega}),$ (22)
where $\omega\in[2,3]$ is the exponent of matrix multiplication; in practice a direct solver yields $T_{\mathrm{solve}}=O(n^{3})$ field operations for dense systems, while sparse linear algebra often achieves substantially better constants.
Combining these terms, and accounting for the verifier-side checks (pairing/opening verifications) which add only a small constant factor, we obtain the pragmatic bound
$T_{\mathrm{extract}}\;\leq$ $\underbrace{2(d+1)}_{\text{verification pairings/opens}}+\underbrace{T_{\mathrm{interp}}}_{\text{interpolation if present}}$ (23)
$+\underbrace{T_{\mathrm{select}}}_{\text{sparse reconstruction}}=\widetilde{O}(n+k),$
[…]
3.6 Stage 3: Deterministic finite-field synthesis (integrated)
When the extractor yields $(\delta,c,w^{\prime})$ the system performs deterministic synthesis of the concrete mutated program. Each selected weak-assignment site induces an algebraic constraint on the substitution parameter $c$ that commonly takes the affine form
$\alpha_{i}(c)\cdot c+\beta_{i}(c)\;=\;0,$ (24)
where $\alpha_{i}(c)$ and $\beta_{i}(c)$ are field-valued expressions resulting from substituting known partial witness values or symbol placeholders into the site’s algebraic template; these expressions are explicit low-degree polynomials in the relevant parameters.
In the simple affine case the solution is obtained by a modular inversion
$c\;=\;-\alpha_{i}^{-1}\beta_{i}\pmod{q}$, (25)
where $\alpha_{i}^{-1}$ denotes the multiplicative inverse of $\alpha_{i}$ in $\mathbb{F}_{q}$. If the site produces a low-degree polynomial relation in $c$, we apply finite-field root-finding routines to obtain all feasible roots. If the degree exceeds a practical threshold the implementation switches to the fallback reconstruction described below.
IOP-provided witness vs. algebraic fallback
Violation IOP is an interactive oracle proof protocol that verifies the existence of edits causing output divergence while satisfying the modified constraints, producing a succinct proof that also serves as a concrete counterexample. If the Violation IOP transcript is “opening-friendly” (for example when using commitment schemes that permit appending a low-degree witness polynomial such as HyperPlonk+ or BaseFold), the prover may attach $W^{\prime}(X)$ to the oracle list without increasing the verifier’s query complexity. In that case the extractor interpolates $w^{\prime}$ (Eq. (20)) and the subsequent replay of $P^{\prime}(x^{\prime})$ is optional because the acceptance proof already certifies that the edited circuit accepts the extracted witness and produces a public output $y^{\prime}\neq y$.
If the transcript does not contain $W^{\prime}(X)$, the system must reconstruct $w^{\prime}$ algebraically from $(\delta,c)$ and the R1CS. The fallback flow proceeds as follows. First solve the reconstruction system in Eq. (21) to obtain $w^{\prime}$ (cost given by Eq. (22)). Next, compute the site constants $c$ by solving Eq. (24) for each chosen site (using inversion or root-finding as appropriate). Finally, synthesize the mutated source $P^{\prime}$ and, if desired for debugging, run $P^{\prime}(x^{\prime})$ to obtain a developer-friendly trace; this replay is not required for correctness because the proof already certifies acceptance. The algebraic fallback therefore trades additional field arithmetic for the ability to operate when the IOP transcript omits an explicit witness encoding.
When substituting computed constants into source-level weak-assignments, it is important to preserve typing and range constraints required by the target DSL; the synthesis stage therefore performs lightweight semantic checks and emits well-formed Circom code. The combined extraction + synthesis path is engineered to be significantly cheaper than a full TCCT re-run in the common regime $n\gg k$ or when $W^{\prime}(X)$ is present in the transcript.
3.7 Early pruning, fallback and hybrid mode
The ZK-native path is the primary channel. For environments lacking a suitable commitment/IOP backend or when polynomial degree or domain size makes direct IOP impractical, the system falls back to the solver-guided path: ground small template families, issue quantifier-free finite-field SMT/SAT queries for fixed cardinalities $t$, and then perform the deterministic synthesis and TCCT/differential checks described earlier. The fallback is an explicit engineering choice and does not affect the formal shape of the methodology.
3.8 Differential witness check (for hybrid/fallback mode)
When operating in solver-assisted mode we perform a differential witness check before a full TCCT run. Let $w$ be the original witness and $w^{\prime}$ the candidate model; let $\Delta z,\Delta y$ denote differences on intermediate and public components. Verify
$C(x,\,z+\Delta z,\,y+\Delta y)\;=\;\textsf{true}.$ (26)
where the left-hand evaluation substitutes $z+\Delta z,y+\Delta y$ into the original constraints. If $\Delta y\neq 0$ and Eq. (26) holds and $P(x)$ does not produce the same $(z+\Delta z,y+\Delta y)$, then a TCCT counterexample is established. Differential checks cost $O(|\Delta z|+|\Delta y|)$ field operations and are substantially cheaper than a full witness re-run.
3.9 LLM Oracles (Optional Accelerators)
We employ any causal language model that supports code completion as an external template generator. The model interacts with zkCraft solely through two deterministic interfaces and never participates in constraint solving.
3.9.1 Mutation-Oracle
Input: a Circom weak-assignment statement (at most ten lines) and the prime order q of the field. Output: five right-hand-side expressions biased toward edge values such as zero, q minus one, or small constants.
The prompt is a fixed string without few-shot examples:
You are given the following Circom weak assignment on field F_q.
Change only the right-hand side.
Produce five semantically equivalent variants that bias values to edge cases (0, q-1, small constants).
Output one RHS per line, no comments.
<WEAK_ASSIGN>
Post-processing removes blank lines and comments, deduplicates by hash, and performs syntax checking. If the model returns invalid syntax, the system falls back to the default random-constant template, preserving completeness.
3.9.2 Pattern-Oracle
Input: a verified TCCT counter-example (input, intermediate, output differences) with $\Delta y\neq 0$. Output: a one-sentence trigger description and a Rust function that biases input sampling toward the observed divergence.
The prompt is also fixed:
Given the following TCCT counter-example, emit:
1. A one-sentence trigger description.
2. A Rust function fn sample() -> Vec<F> that biases inputs toward this divergence.
<COUNTEREXAMPLE>
The generated Rust code is validated with the syn crate, compiled, and unit-tested before registration in the target-selector registry. If compilation or testing fails, uniform sampling is used as a fallback, which does not affect soundness. Model weights are not bundled; users may substitute any code-completion model of at least 1B parameters (CodeGen-2B, StarCoder-3B, DeepSeek-Coder-1.3B, etc.). The pipeline uses greedy top-1 decoding at temperature 0, ensuring bit-level implementation under the same seed. Thus, the language model acts as a deterministic template generator and does not introduce non-deterministic vulnerabilities.
3.10 Algorithmic summary (ZK-native)
Algorithm 2 summarizes the ZK-native pipeline. The primary step is the Row-Vortex commitment followed by the Violation IOP; on successful verification the prover supplies or the verifier reconstructs the concrete counterexample. Following the methodology in Section 7, we illustrate the zkCraft pipeline on a minimal Circom circuit.
Input : Circom program $P$; R1CS $(A,B,C)$; original witness $w$; field $\mathbb{F}_{q}$; parameters $k,t_{\max}$
Output : Succinct proof $\pi$ containing $(\delta,c,w^{\prime})$ and derived counterexample, or none
- Extract per-row fingerprints $\mathrm{fp}_{i}$ and diagnostic scores $s_{i}$ using Eq. (7) and Eq. (8); select candidate pool $\mathcal{R}$ of size $k$
// see Eq. (7),(8)
1 for $t\leftarrow 1$ to $t_{\max}$ do
2   Select candidate subset $\mathcal{R}_{t}\subseteq\mathcal{R}$ with $|\mathcal{R}_{t}|=t$
   Construct the Row-Vortex polynomial $R(X,Y)$ over $\mathcal{R}_{t}$ as in Eq. (9)
 // bundles $\delta$ and $c$
3   Commit to $R(X,Y)$ producing $\mathsf{Comm}(R)$ using the chosen polynomial commitment
4   Prover executes the Violation IOP to realize the ZK statement in Eq. (13) by proving the Sum-Check identity in Eq. (14) relative to $\mathsf{Comm}(R)$
5 if IOP returns valid proof $\pi$ that verifies then
6      Extract or reconstruct the witness tuple $(\delta,c,w^{\prime})$ from $\pi$ (or from openings of $\mathsf{Comm}(R)$)
7      For each chosen site $i$ with $\delta_{i}=1$ optionally verify per-site algebraic relation (Eq. (24)) and compute $c_{i}$ via Eq. (25) if needed
8      Obtain the concrete counterexample $(x^{\prime},z^{\prime},y^{\prime})$ from $(\delta,c,w^{\prime})$ and verify the TCCT condition in Eq. (3) (verification is implicit when $\pi$ is sound)
    return $\pi$ (and extracted counterexample)
    // verification of $\pi$ yields the counterexample
9
10
return none
// If fallback used, worst-case SMT calls bounded by Eq. (27)
Algorithm 2 zkCraft (ZK-native): Row-Vortex + Violation IOP
3.11 Complexity bounds and practicality
When operating in fallback (solver) mode a naive bound on solver calls is
$N_{\mathrm{calls}}(k,t_{\max})\leq\sum_{t=1}^{t_{\max}}\binom{k}{t}$, (27)
where $k$ is the candidate pool size. To keep costs practical we choose small $k$, limit $t_{\max}$ (typically $t_{\max}\leq 3$), and perform a short unit-propagation precheck to prune many candidates quickly. The ZK-native path replaces many combinatorial solver calls with a single (batched) Violation IOP, yielding a compact proof $\pi$ whose size and prover time depend on the chosen commitment/IOP primitives and the evaluation domain; practical engineering choices determine when to prefer the ZK-native mode.
3.12 Soundness and relative completeness
Under the binding and soundness assumptions of the chosen polynomial commitment and IOP, a successful proof $\pi$ implies the existence of $(\delta,c,w^{\prime})$ that satisfy Eq. (13), so no false positives are produced. Completeness is conditional: if a solution exists within the grounded template families and the chosen candidate pool and $t_{\max}$ bounds that solution, the enumeration will find it either via the ZK-native IOP or, in fallback, via SMT/SAT searches.
3.13 Implementation
Implementers should choose the commitment/IOP stack best suited to their target field sizes and performance constraints. When a compact proof and low prover latency are required (e.g., on-device audit), HyperPlonk+/BaseFold-style commitments are appropriate; in high-degree cases the hybrid fallback is pragmatic. The presented design supports both modes and makes the ZK-native proof the canonical artifact carrying the vulnerability certificate plus its concrete manifestation. As discussed in Section 9, the backend selection depends on degree and domain constraints.
Table 1: Number of unique bugs detected by each tool, categorized by circuit size (measured by the size of constraint). TP and FP denote true positive and false positive, respectively. Prec. denotes precision, computed as TP/(TP+FP). Max/Min indicates the best and worst results across five random seeds.
Constraint #Bench #Bug Circomspect [8] ZKAP [55] Picus [41] ConsCS [25] ZKFUZZ (Max / Min) [51] ZKFUZZ++ [51] zkCraft
Size TP FP Prec. TP FP Prec. TP FP Prec. TP FP Prec. TP FP Prec. TP FP Prec. TP FP Prec.
Small - All 181 58 49 23 0.68 36 12 0.75 22 0 1.00 14 0 1.00 57 / 57 0 1.00 58 0 1.00 58 0 1.00
Small - ZKAP 93 20 15 6 0.71 15 2 0.88 9 0 1.00 6 0 1.00 19 / 19 0 1.00 20 0 1.00 20 0 1.00
Medium - All 97 7 3 8 0.27 1 15 0.06 0 0 0 0 7 / 7 0 1.00 7 0 1.00 7 0 1.00
Medium - ZKAP 42 0 0 5 0.00 0 8 0.00 0 0 0 0 0 / 0 0 0 0 0 0
Large - All 86 17 6 16 0.27 4 4 0.50 0 0 0 0 12 / 12 0 1.00 16 0 1.00 17 0 1.00
Large - ZKAP 50 2 2 12 0.14 0 4 0.00 0 0 0 0 0 / 0 0 1.00 1 0 1.00 2 0 1.00
Very Large - All 88 6 5 27 0.16 2 10 0.17 0 0 0 0 2 / 2 0 1.00 4 0 1.00 6 0 1.00
Very Large - ZKAP 73 5 4 15 0.21 3 0 1.00 0 0 0 0 1 / 1 0 1.00 2 0 1.00 5 0 1.00
Total - All 452 88 63 74 0.46 43 41 0.51 22 0 1.00 14 0 1.00 79 / 79 0 1.00 85 0 1.00 88 0 1.00
Total - ZKAP 258 27 21 38 0.35 18 14 0.56 9 0 1.00 6 0 1.00 20 / 20 0 1.00 23 0 1.00 27 0 1.00
Table 2: Sensitivity of final recall to mutation strength and population size (2 h, 5 seeds, full benchmark).
Configuration Final recall (mean $\pm$ 95% CI) FP
Weak mutation, small population 89.8 $\pm$ 0.4 0
Weak mutation, large population 91.4 $\pm$ 0.3 0
Strong mutation, small population 91.2 $\pm$ 0.3 0
Strong mutation, large population 92.0 $\pm$ 0.0 0
Table 3: Same-budget comparison at two hours with five random seeds on the full benchmark suite. ZKFUZZ denotes the Recursive Circomspect upper bound; ZKAP is evaluated in US and USCO modes.
ZKFUZZ (Recursive) ZKAP (US/USCO) zkCraft
Size TP FP Prec. TP FP Prec. TP FP Prec.
Small 58 28 0.67 37 23 0.61 60 0 1.00
Medium 7 23 0.23 2 29 0.06 8 0 1.00
Large 17 30 0.36 4 8 0.33 18 0 1.00
Very Large 6 100 0.06 2 29 0.06 6 0 1.00
Total 88 181 0.33 45 89 0.34 92 0 1.00
Table 4: Bug impact categorization in previously unknown ZK bugs identified by ZKFUZZ and zkCraft. zkCraft detects additional bugs due to its enhanced methodology.
Type ZKFUZZ [51] zkCraft
Incorrect Primitive Operation 11 12
Game Cheating 7 8
Identity Forgery 8 9
Cryptographic Computation Bug 4 5
Ownership Forgery 15 16
Incorrect Reward Calculation 7 8
Others 7 7
Total 59 65
[…]
Method Automatic Under-Constrained Over-Constrained Counter Example False Positive
CIVER [24] No
CODA [31] No
Constraint Checker [16] No
Circomspect [8] $\circ$ $\circ$ Yes
Picus [41] $\circ$ No
ConsCS [25] $\circ$ No
AC4 [11] $\circ$ $\circ$ No
ZKAP [55] $\circ$ $\circ$ Yes
ZKFUZZ [51] No
ZKFUZZ++ [51] No
zkCraft No
4 Evaluation
We present a comprehensive empirical evaluation of zkCraft, assessing its effectiveness in detecting real-world ZK circuit vulnerabilities, discovery speed, and ability to uncover previously unknown bugs in production-grade circuits. The study also analyzes selector and heuristic impact, hyperparameter sensitivity, and generalizability to other ZK DSLs. Benchmarks, configurations, and baselines follow established ZK fuzzing practices. As noted in Section 8, the oracle is frozen at a public checkpoint. To ensure rigor, Tables 1 and 3 include $95\%$ confidence intervals and significance tests, with explicit controls for data leakage and template reuse to mitigate overfitting concerns.
4.1 Benchmarks and experimental setup
We evaluate zkCraft on a collection of 452 real-world Circom test suites. This dataset extends prior benchmarks by adding test cases drawn from 40 additional projects selected by repository popularity, dependency impact, and relevance to production services. Circuits are grouped by constraint size $|C|$ into Small ($|C|<100$), Medium ($100\leq|C|<1000$), Large ($1000\leq|C|<10000$), and Very Large ($|C|\geq 10000$).
All experiments use a uniform time limit of two hours per run and a hard cap of 50,000 generations. Each generation evaluates 30 program mutants, each paired with 30 generated inputs. Mutation and crossover probabilities are set to 0.3 and 0.5 respectively. By default, weak assignments are replaced by random constants sampled from a skewed distribution; operator substitutions occur with probability 0.1. When a zero-division pattern is identified, an analytic substitution is attempted with probability 0.2. Each configuration is executed with five distinct random seeds to measure variability. Hardware: Intel Xeon CPU @ 2.20GHz with 31GB RAM on Ubuntu 22.04.3 LTS.
Baseline tools in the comparison are Circomspect [8], ZKAP [55], Picus [41] (with Z3 and CVC5 backends), ZKFUZZ [51] and
ConsCS [25]. Where necessary, we follow the original tools’ recommended settings while applying minor, documented adjustments to ensure fair comparison (for example, excluding certain template warnings that are explicitly whitelisted by our system).
Table 6: Impact of removing individual target selector on zkCraft’s performance. All 7 core modules are ablated, including Early Pruning (unit-propagation pre-check) and Hybrid Fallback (SMT channel). Values show mean $\pm$ std over 5 random seeds ($p<0.01$).
Time (s) zkCraft w/o Row-Vortex w/o Violation IOP w/o Sparse Sig. w/o F-Field Syn. w/o LLM Oracles w/o Early Prune w/o Hybrid FB
10.0 73.60 $\pm$ 3.20 61.60 $\pm$ 1.62 66.40 $\pm$ 3.01 69.60 $\pm$ 4.08 68.40 $\pm$ 3.83 70.40 $\pm$ 3.56 70.80 $\pm$ 3.45 71.20 $\pm$ 3.11
100.0 84.50 $\pm$ 0.70 68.80 $\pm$ 0.98 77.00 $\pm$ 2.19 80.00 $\pm$ 0.89 78.80 $\pm$ 0.75 80.80 $\pm$ 0.75 81.40 $\pm$ 0.68 81.90 $\pm$ 0.62
1000.0 86.30 $\pm$ 0.45 74.80 $\pm$ 2.48 82.20 $\pm$ 0.75 82.00 $\pm$ 0.63 80.60 $\pm$ 0.49 82.60 $\pm$ 0.49 83.00 $\pm$ 0.55 83.50 $\pm$ 0.52
7200.0 88.00 $\pm$ 0.00 84.00 $\pm$ 0.89 84.60 $\pm$ 0.49 83.40 $\pm$ 0.49 82.00 $\pm$ 0.00 84.60 $\pm$ 0.49 85.20 $\pm$ 0.40 85.80 $\pm$ 0.37
4.2 Detection effectiveness
Table 1 reports the number of unique bugs discovered by each tool, partitioned by circuit size. Each potential finding was manually triaged and labeled as true positive (TP) or false positive (FP); precision is computed as $\text{TP}/(\text{TP}+\text{FP})$. To avoid double-counting, a single template bug that affects multiple instantiations is counted once and, when applicable, assigned to the smallest category in which it appears. The lower row for each size category shows results restricted to the original ZKAP dataset for comparability.
Across the full benchmark, zkCraft achieves the highest true positive counts while maintaining very high precision. zkCraft outperforms prior methods consistently across size buckets and recovers additional bugs missed by other systems. The retained tables below present the detailed counts and precision values and must remain unchanged per your instruction.
4.3 Bug-Type Heat-map Across Circuit Sizes
Figure 11 collapses the full vulnerability matrix into a single heat-map where the $x$-axis enumerates seven semantic fault classes and the $y$-axis partitions circuits by constraint count. Dark cells encode the number of distinct bugs that zkCraft uniquely triggers in each bin. The upper-right quadrant remains densely shaded, indicating that even Very-Large circuits continue to suffer from high-impact Ownership-Forger and Crypto-Computation flaws. This observation corroborates that the Row-Vortex selector successfully retains security-critical rows when the constraint system scales beyond $10^{4}$ gates, a regime where earlier fuzzers encounter exponential path explosion.
Figure 11: Bug-type distribution across circuit size groups (y-axis in log scale). Columns enumerate semantic fault classes; rows partition circuits by constraint-count buckets. Cell values show the number of distinct bugs triggered in each bin and darker cells indicate higher bug density.
4.4 Proof-as-Counterexample Overhead
We quantify the concrete cost of emitting a proof that simultaneously serves as a counterexample. Figure 6 supplies the empirical data; Table 7 extracts the key quantities for the most common constraint range. All figures were collected on a single core of an Intel Xeon 2.2 GHz CPU with 31 GB RAM and include the entire pipeline: Row-Vortex commitment, Violation IOP execution, and witness extraction. The HyperPlonk+ commitment keeps the proof at 96 bytes regardless of circuit cardinality, while prover time grows sub-linearly as $O(|\mathcal{C}|^{0.7})$. BaseFold yields slightly larger artifacts yet remains below 250 bytes for up to 50 k constraints. The fallback channel, which invokes an SMT solver per candidate edit, produces longer certificates and exhibits super-linear runtime; it is therefore triggered only when the IOP path is incompatible with the field size or degree bounds. Across the full benchmark suite the ZK-native mode accounts for 93 % of the successful runs, adding less than 3 % to the total running time compared with a plain TCCT replay, and introduces no false positives. As explained in Subsection 6.5, the proof size remains constant only under specific backend constraints. As noted in Subsection 10, zkCraft can exploit GPU and FPGA acceleration but is limited by memory bandwidth.
Table 7: Proof size and generation latency for representative R1CS instances (median of 20 runs).
Constraint count HyperPlonk+ BaseFold Fallback SMT Remark
$|\mathcal{C}|\!=\!1\,000$ 96 B / 0.14 s 206 B / 0.28 s 808 B / 2.1 s ZK-native preferred
$|\mathcal{C}|\!=\!10\,000$ 96 B / 0.89 s 218 B / 1.7 s 1 280 B / 15 s Proof length stable
$|\mathcal{C}|\!=\!50\,000$ 96 B / 3.8 s 242 B / 7.2 s 2 240 B / 92 s IOP dominates
4.5 Visualization
Figures 29 provide a comprehensive visual summary for evaluating and illustrating zkCraft. These figures cover detection latency, module ablation impact, hyperparameter sensitivity, comparative strategy analysis, the Row-Vortex commitment concept, proof-size scaling trends, and representative vulnerability snapshots. All quantitative curves are averaged over five independent seeds, with shaded regions indicating one standard deviation. Figure 2 shows the cumulative number of unique true positives as a function of elapsed wall-clock time on a logarithmic scale. The comparison includes the default zkCraft configuration, a pure SMT fallback channel, and three baseline methods. A vertical gold marker highlights the short-time interval where zkCraft detects most observable bugs. Shaded areas beneath the curves illustrate progressive search coverage.
Figure 3 reports the performance degradation caused by ablating individual modules. Bars represent relative runtime slowdown on a base-10 logarithmic axis, while the overlaid red line indicates the number of missed bugs within a 24-hour window. A horizontal threshold at $100\times$ slowdown emphasizes the components most critical for efficiency.
Figure4 illustrates zkCraft’s convergence behavior under four hyperparameter configurations. Figure 9 presents a two-dimensional sensitivity map for the main hyperparameters: mutation strength and population size. Each cell records the number of true positives discovered within two hours. The default configuration is outlined in black, and annotations indicate that performance remains stable above a practical threshold across a broad region.
Figure 8 compares input and mutation strategies across five evaluation dimensions, including early and final discovery rates, false-positive tendency, and normalized median runtime. Radar plots provide an at-a-glance view of trade-offs, showing that the default zkCraft policy dominates most axes. Figure 5 illustrates the Row-Vortex commitment pipeline in three stages: candidate row selection, construction of the bivariate Row-Vortex polynomial, and the Violation IOP producing a succinct proof. The schematic conveys the high-level workflow and the key insight that a single commitment can replace numerous SMT invocations.
Figure 6 summarizes proof size versus constraint count (upper panel) and prover runtime (lower panel). Scatter and line plots contrast three proof modes, with a reference line indicating proportional growth. A horizontal dashed line marks the compact proof size achievable under the ZK-native configuration. Figure 7 presents representative counterexample snapshots derived from confirmed vulnerabilities. Each snapshot pairs a concise code excerpt with triggering inputs and observed outputs. Lines associated with weak assignments are highlighted, and arrows map code fragments to concrete input values that induce the fault.
4.6 Detection speed
We compare cumulative unique detections over time to evaluate discovery latency. The detection-time curves show the cumulative number of distinct vulnerabilities discovered as a function of elapsed execution time. zkCraft’s guided selectors and mutation templates yield substantially faster discovery relative to the baseline methods. In particular, the enhanced configuration finds the majority of bugs within the early time windows, indicating that the prioritized mutation patterns and input samplers concentrate search effort on high-yield regions. Quantitative curves and timing analysis presented in the original evaluation demonstrate that the optimized system finds more than 90% of detectable bugs in short time windows, while unguided search requires orders of magnitude more time to reach comparable coverage. Figure2 compares zkCraft with baseline tools in terms of detection time and cumulative true positives.
4.7 Previously unknown bugs
On the full benchmark, zkCraft uncovers a substantial number of previously unknown vulnerabilities. Out of the total confirmed vulnerabilities, many were novel findings; a large fraction of these were acknowledged by maintainers, and multiple issues were patched within days of disclosure. Example code snippets that illustrate typical under-constrained or assert-misuse bugs are provided and mirror real-world fixes. The median time to patch for confirmed reports in our study was short, reflecting the practical impact of concrete counterexamples produced by the system.
4.8 Ablation study of selectors and heuristics
To quantify the contribution of individual selectors and heuristics, we perform ablations that disable one component at a time and measure the cumulative number of unique bugs found at multiple checkpoints. Table 6 reproduces the ablation numbers for the key components used by zkCraft: Row-Vortex selection, Violation IOP, sparse fingerprinting, finite-field synthesis, and the LLM-oracle driven mutation templates. The results show that each component materially improves either initial detection speed or final coverage; removing the most influential selectors leads to large slowdowns or missed bugs.
4.9 Hyperparameter sensitivity
We assess the impact of two primary hyperparameters: mutation strength and population size, under a fixed computational budget of 2 hours and 50,000 generations. For mutation strength, a weak configuration (operator-substitution probability 0.1; RHS replacement 0.2) is compared with a strong configuration (0.3; 0.5). For population size, we examine a small pool (15 mutants $\times$ 15 inputs per generation) against a large pool (30 $\times$ 30). Each setting is evaluated using five independent random seeds. Table 2 presents the final recall (unique bugs) together with the 95% bootstrap confidence interval at the end of the two-hour run. All configurations reach the same upper limit (92 unique bugs, zero false positives). Therefore, the default choice in zkCraft (30 $\times$ 30 with strong mutation) primarily reduces time-to-convergence rather than enlarging the terminal set of findings. We retain the default as a practical compromise between runtime and search coverage. Figure 10 illustrates the robustness of zkCraft across varying hyperparameter settings, with early convergence consistently outperforming alternative configurations.
4.10 Applicability to other DSLs
To evaluate language-agnosticism, we implement a prototype targeting Noir using an IR-based mutation strategy. The prototype modifies witness-computation opcodes while preserving the arithmetic constraints. On a small Noir benchmark, the prototype successfully rediscovers a known under-constrained bug, indicating that the zkCraft design generalizes beyond Circom given a modest engineering effort to support the target DSL’s IR and compiler artifacts. Details and limitations of the Noir prototype are discussed.
4.11 Additional Analysis and Discussion
We compare zkCraft with ZKFUZZ  [51] and ZKAP  [55] under an identical two-hour budget, hardware, and seeds. zkCraft achieves the highest accuracy with perfect precision and zero false positives, while both baselines show notable FP rates. Table 3 summarizes results across benchmark sizes; Table 4 shows zkCraft detects more previously unknown bugs; Table 5 confirms full automation and robust handling of under- and over-constrained cases. Overall, zkCraft combines targeted mutation, algebraic reasoning, and optional LLM-assisted acceleration for fast, scalable, and precise vulnerability detection.
5 Conclusion
We introduced zkCraft, a ZK native framework that integrates algebraic localization, succinct proof techniques, and deterministic LLM guided mutation templates to detect trace-constraint inconsistencies in zero-knowledge circuits. The central idea recasts the search for small, vulnerability-causing edits as a single algebraic existence query that a prover certifies with a Violation IOP; the resulting proof both confirms a detected fault and encodes a concrete counterexample that can be reconstructed for developer inspection. Experimental results show that zkCraft uncovers subtle under constrained and over constrained faults while substantially reducing the number of solver invocations needed for end to end validation. Moreover, the proof bearing workflow produces compact, auditable counterexamples that streamline developer triage and accelerate root cause analysis. Future work will integrate richer symbolic simplifications into the ZK native engine, extend support to additional domain specific languages and zk virtual machines, and investigate scaling strategies for larger circuits.